OpenStudy (anonymous):

How much work must be done to increase the speed of an electron (a) from 0.18c to 0.19c and (b) from 0.98c to 0.99c? Note that the speed increase is 0.01c in both cases.

OpenStudy (anonymous):

Well - do you know about Lorentz transformation and calculating the Lorentz factor? Also, before you start doing any calculations.. do you expect the work to be different or the same for both cases? How? Why?

OpenStudy (anonymous):

For part (a), it's 1.0 ke V and for part (b), it's 1.1 Me V. Now tell me how to find them. I know how to find the Lorentz factor.

OpenStudy (anonymous):

Again: Do you expect the work to be different or the same for both cases? How? Why? And: NO, I'm not just simply gonna tell you how to find them - I am however trying to help and guide you there. Now how about you try to explain to me/us what you think is happening in this example?

OpenStudy (anonymous):

I got 1*10^-16 but what's the unit for that?

OpenStudy (anonymous):

Are you friggin kidding? You neither tell me what you did, or how you did it .. or why.. and you refuse to answer questions yourself. Yet somehow, you think I would be able to magically know what you are talking about..

OpenStudy (anonymous):

I'm not kidding you. Here's what I did. If you don't want to help, then simply leave. W=(9.11x10^-31)(((0.19)^2-(0.18)^2))/2)=1.69*10^-33 J=1.05*10^-14 eV=1.05*10^-17 keV What I don't understand is the units in the textbook. Maybe the answers in the book was wrong but aren't the units should be the same in both parts (a) and (b)?